Circular No. 8540

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

POSSIBLE NOVA IN AQUILA G. Pojmanski reports the ASAS detection (cf. IAUC 8495, 8539) of

a variable star located at α = 19^{h}05^{m}12^{s}, δ = +5^{o}14′.2 (equinox 2000.0), providing the following V magnitudes: June 3.318 UT, [14; 9.240, 11.05; 10.226, 10.5:; 10.234, 10.8: (the June 10 frames taken through heavy cloud cover). A. Oksanen, Jyväskylä, Finland, reports the variable star at V = 10.84 (noting the object as being very red on images taken in various passbands) on June 10.36 from a remote observation with the 0.30-m Mewlon telescope in New Mexico, providing the following position end figures: 12^{s}.50, 12′′.0. Oksanen adds that a USNO-B1.0 star (red mag 18.1) has the following position end figures: 12^{s}.186, 12′′.25.

SUPERNOVA 2005ch Further to IAUC 8539, E. J. Christensen reports that SN 2005ch ap-

peared at V = 17.3–17.4 on prediscovery images obtained on May 31.27– 31.29 UT in the course of the Catalina Sky Survey. Nothing is visible at the location of the new object on Catalina Sky Survey images obtained on May 19.14 (limiting mag V ≈ 19.5). R. A. Tucker, Tucson, AZ, reports the following unfiltered CCD magnitudes obtained with the 0.35-m MOTESS telescopes: May 30.176, 17.5; June 1.170, 17.1; 3.192, 16.5; 4.189, 16.4; 5.186, 16.2; 6.183, 16.1; 7.167, 16.1; 8.165, 16.0; 9.162, 15.9.

VARIABLE STAR IN NORMA Further to IAUC 8539, L. A. G. Monard, Pretoria, S. Africa, reports

the following precise position from a frame taken by himself on June 9.87 UT (when the star was at R_{c} = 12.55 and V = 12.75): α = 16^{h}00^{m}47^{s}.43, δ = −48^{o}46′07′′.6 (equinox 2000.0; uncertainty ± 0′′.25). He adds that this lies close to a USNO-A2.0 star with position end figures 47^{s}.475, 05′′.43 (blue mag 17.9, red mag 15.7). Six hours of time-series unfiltered photometry showed an initial large hump in the light curve, which gradually was replaced by smaller humps and a general fading trend.

COMET C/2005 K2 (LINEAR) This comet appears to have had an outburst; total visual magnitude es-

timates: June 6.48 UT, 11.7 (S. Yoshida, Gunma, Japan, 0.40-m reflector); 7.97, 10.3 (N. Biver, Ablis, France, 0.41-m refl.); 9.08, 9.8 (J. J. Gonzalez, Leon, Spain, 25×100 binoculars); 9.95, 8.9 (Gonzalez, 11×80 binoculars).

2005 June 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green